If a cup of ice cream is selected. what is the probability that the cholesterol content will be more than 670mg?


1
Expert's answer
2022-04-28T13:23:07-0400

Given "\mu=660\ mg, \sigma=35\ mg, n=10"

"P(X>670)=1-P(X\leq670)""=1-P(Z\leq\dfrac{670-660}{35/\sqrt{10}})""\approx1-P(Z\leq0.9035)\approx0.1831"

The probability that the mean of the sample will be larger than 670 mg is 0.1831.